# Identifying biological perturbation targets through causal differential networks

### v2

@@ Abstract
Identifying variables responsible for changes to a biological system enables applications in drug target discovery and cell engineering. Given a pair of observational and interventional datasets, the goal is to isolate the subset of observed variables that were the targets of the intervention. Directly applying causal discovery algorithms is challenging: the data may contain thousands of variables with as few as tens of samples per intervention, and biological systems do not adhere to classical causality assumptions. We propose a causality-inspired approach to address this practical setting. First, we infer noisy causal graphs from the observational and interventional data. Then, we learn to map the differences between these graphs, along with additional statistical features, to sets of variables that were intervened upon. Both modules are jointly trained in a supervised framework, on simulated and real data that reflect the nature of biological interventions. This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets. We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.
@@ 1 Introduction
Machine learning methods have primarily been designed to infer the effects of perturbations on cells, with the goal of generalizing to unseen perturbations (Roohani et al., 2023), or unseen cell distributions (Lotfollahi et al., 2023; Bunne et al., 2023). In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2023; Zhang et al., 2023a). However, the number of candidate sets scales exponentially with the number of variables under consideration, and there are limited training data for combinatorial perturbations ($<150$ pairs in Norman et al. (2019)). Thus, this option may be impractical for larger search spaces.
@@
In this paper, we propose causal differential networks (Cdn): a causality-inspired approach to identify variables that drive desired shifts in cell state, while estimating their mechanistic structure directly from data. Given a pair of observational and interventional datasets, we train a causal structure learner to predict causal graphs that could have generated each dataset. The pair of graphs is input to an attention-based classifier, which predicts whether each variable was subject to intervention. To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2022; Petersen et al., 2023), over thousands of synthetic or real datasets.
We evaluate Cdn on real transcriptomic data and synthetic settings. Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches) on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024), without using any external knowledge. Furthermore, Cdn generalizes with minimal performance drop to unseen cell lines, which have different supports (genes), causal mechanisms (gene regulatory networks), and data distributions. Finally, on synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
@@ 2.2 Modeling perturbations
Each experiment is conducted in a single cell line, which represents a distinct set of variables $X$, distribution $P_{X}$, and graph $G$. These three objects may share similarities across cell lines, as certain genes or pathways are essential for survival. However, the empirical discrepancy is quite large between experiments. For example, Nadig et al. (2024) reports that the median correlation in log-fold change of genes under the same perturbation, across two experiments in the same cell line, by the same lab, was 0.16.
There are two directions for modeling perturbations. To reduce experimental costs, which scale with the number of cell lines and perturbations, one option is to infer the post-intervention distribution of cells. That is, given $D_{\text{obs}}\sim P_{X}$ and $I$, the goal is to infer $\tilde{P}_{X}$. Works such as scGen (Lotfollahi et al., 2019), CPA (Lotfollahi et al., 2023), ChemCPA (Hetzel et al., 2022), and CellOT (Bunne et al., 2023) aim to generalize to new settings $P_{X}$, e.g. unseen cell lines or patients. Alternatively, models like GEARS (Roohani et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) are designed to generalize to unseen perturbations. These approaches do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2023). For example, given an oracle $\phi:D_{\text{obs}},I\to\tilde{P}$, the $I^{*}$ associated with $\tilde{P}^{*}$ is
@@
for some dissimilarity $d$. However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting potential barriers to this approach.
Perturbation datasets can also be used to train models for predicting targets of interest, as is the focus of this work. A common strategy is to relate observed changes (e.g. disease, gene levels) to mechanistic explanations (e.g. pathways), in which the targets are involved. To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2024). However, these information are collected from highly inhomogenous sources (Ashburner et al., 2000), and it is difficult to quantify their correctness and completeness in each new setting. In contrast to these approaches, we choose to infer the mechanistic structure directly from data.
@@ 2.3 Causal discovery
If data from multiple environments are available, they can provide additional identifiability regarding the system (Jaber et al., 2020; Zhang et al., 2023b). Differences between these environments, e.g. intervention targets $I$, are either given as input or part of the inference task. As in the observational case, discovery algorithms can be categorized into discrete search (Hauser & Bühlmann, 2012; Ghassami et al., 2018; Ke et al., 2019; Huang et al., 2020; Mooij et al., 2020) or continuous optimization (Brouillard et al., 2020; Hägele et al., 2023). Additionally, Varici et al. (2022) and Yang et al. (2024) frame intervention target prediction as its own task, similar to this paper. However, the former strictly assumes linearity, while the latter requires at least as many environments as $\|I\|$ (we only consider before and after perturbation). Finally, Dci (Belyaeva et al., 2021) infers edge-level differences between linear Gaussian models, rather than identifying the full causal structure. Our work is also motivated by the idea that differences between summary statistics (e.g. regression coefficients in Dci) are informative of changes between systems. Thus, Cdn can be viewed as implementing similar operations within an end-to-end, supervised framework.
@@ 2.4 Amortized causal inference
While the majority of causal discovery algorithms must be fit or run from scratch for each data distribution, amortized causal discovery algorithms have addressed this task as a supervised machine learning problem. Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2022; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2024). Compared to traditional approaches, these algorithms trade explicit assumptions for implicit characteristics of the data simulation process (Montagna et al., 2024). These models are also fast and robust to noise. In this work, we simulate interventions that reflect biological perturbations, and our model architecture builds upon that of Wu et al. (2024), as it scales easily to hundreds of variables.
@@ 3.1 Causal structure learner
The causal structure learner takes as input dataset $D$ and outputs the predicted $G$. In principle, this module can be implemented with any data-efficient causal discovery model. Here, we based the architecture on Sea (Wu et al., 2024), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
Input featurization  We represent dataset $D$ through two types of statistics. First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$, like pairwise correlation, over all variables. Second, we obtain estimates of local causal structure by running classical causal discovery algorithms like FCI (Spirtes et al., 1995) over small subsets of variables. We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types (e.g. ,$\leftrightarrow$,-$\rightarrow$ , - , →), and $k,T$ are hyperparameters that trade off accuracy vs. runtime.
@@
where $h\in\mathbb{R}^{\dots\times d}$ denotes an intermediate hidden representation. This design takes after Transformer models for higher dimensional data in other domains, e.g. Axial Transformers for videos (Ho et al., 2020) and MSA Transformers for aligned protein sequences (Rao et al., 2021). A key difference is the graph modality: when making edge or node-level predictions, the model should be equivariant to node labeling. This is accomplished by adding randomly-permuted “positional” embeddings to each node and edge representation. For more details, please see Appendix A.
@@ 3.3 Implementation details
Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995), and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable. We swept over the number of differential network layers (Figure 4) on synthetic data, and we used 3 layers for $h_{\text{cat}}$ and 2 layers for $h_{\text{diff}}$. Following Sea, we adopted hidden dimension $d=64$, the AdamW optimizer (Loshchilov & Hutter, 2019), learning rate 1e-4, batch size 16, and weight decay 1e-5. On the real data, where $N=1000$, we changed to a batch size of 1, decreased the learning rate to 5e-6, and finetuned the models with half precision (FP16). We trained both $h_{\text{cat}}$ and $h_{\text{diff}}$ models on synthetic data, but only finetuned the $h_{\text{diff}}$ architecture on the real data (memory constraint).
@@ 4 Experiments
We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1). For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
@@ 4.1 Biological experiments
Datasets  We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al. (2022) and Nadig et al. (2024); as well as two Sci-Plex (Srivatsan et al., 2020) datasets (chemical perturbations) from McFaline-Figueroa et al. (2024). Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$. In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target (discussion in Appendix D.1).
@@
  • We compute the rank of the true target, normalized by the number of candidate genes. Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
@@
  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates. Recall ranges from 0 to 1, where 1 is best.
@@ A.1 Attention layers
To balance expressivity and efficiency, our model is designed around series of “axial” attention layers, similar to Ho et al. (2020); Rao et al. (2021); Wu et al. (2024). The causal structure learner closely resembles the architecture of Sea (Wu et al., 2024), while the differential network is slightly different (since the input space is different). Briefly, the input to the causal structure learner is a matrix of shape $N\times N\times T\times d$. However, the (dense) global statistic does not span the $T$ dimension, and the local graph estimates are generally sparse. Thus, the causal graph learner separately attends over the $N\times N\times d$ and $K\times T\times d$ components ($K$ is the number of unique edges in sparse format), with “messages” to align the representations between each attention layer. In contrast, the differential network only operates over the dense $N\times(N+1)\times d$ (or $2d$) representation. Thus, it omits the message passing layers and directly stacks dense attention layers.

### v4

@@ Abstract
Identifying variables responsible for changes to a biological system enables applications in drug target discovery and cell engineering. Given a pair of observational and interventional datasets, the goal is to isolate the subset of observed variables that were the targets of the intervention. Directly applying causal discovery algorithms is challenging: the data may contain thousands of variables with as few as tens of samples per intervention, and biological systems do not adhere to classical causality assumptions. We propose a causality-inspired approach to address this practical setting. First, we infer noisy causal graphs from the observational and interventional data. Then, we learn to map the differences between these graphs, along with additional statistical features, to sets of variables that were intervened upon. Both modules are jointly trained in a supervised framework, on simulated and real data that reflect the nature of biological interventions. This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets. We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.111 Code is available at https://github.com/rmwu/cdn
@@ 1 Introduction
Machine learning methods have primarily been designed to infer the effects of perturbations on cells, with the goal of generalizing to unseen perturbations (Roohani et al., 2023), or unseen cell distributions (Lotfollahi et al., 2023; Bunne et al., 2023). In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2024; Zhang et al., 2023a). However, the number of candidate sets scales exponentially with the number of variables under consideration, and there are limited training data for combinatorial perturbations ($<150$ pairs in Norman et al. (2019)). Thus, this option may be impractical for larger search spaces.
@@
In this paper, we propose causal differential networks (Cdn): a causality-inspired approach to identify variables that drive desired shifts in cell state, while estimating their mechanistic structure directly from data. Given a pair of observational and interventional datasets, we train a causal structure learner to predict causal graphs that could have generated each dataset. The pair of graphs is input to an attention-based classifier, which predicts whether each variable was subject to intervention. To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2023; Petersen et al., 2023), over thousands of synthetic or real datasets.
We evaluate Cdn on real transcriptomic data and synthetic settings. Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches), evaluated on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024) without using any external knowledge. Furthermore, Cdn generalizes with minimal performance drop to unseen cell lines, which have different supports (genes), causal mechanisms (gene regulatory networks), and data distributions. On synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
@@ 2.2 Modeling perturbations
Each experiment is conducted in a single cell line, which represents a distinct set of variables $X$, distribution $P_{X}$, and graph $G$. These three objects may share similarities across cell lines, as certain genes or pathways are essential for survival. However, the empirical discrepancy can be quite large between experiments. For example, Nadig et al. (2024) reports that the median correlation in log-fold change of genes under the same perturbation, across two experiments in the same cell line, by the same lab, was 0.16.
Early efforts to map the space of cells measured the distribution of gene levels as a surrogate for cell state (Lamb et al., 2006). Diseases, drugs, or genetic perturbations that induce similar changes in this distribution were thought to act through similar mechanisms. However, it is intractable to map all combinations of starting cell states and perturbations through experiment alone. This bottleneck has led to two directions in the in-silico modeling of perturbations.
To reduce experimental costs, which scale with the number of cell lines and perturbations, one option is to infer the post-intervention distribution of cells. That is, given $D_{\text{obs}}\sim P_{X}$ and $I$, the goal is to infer $\tilde{P}_{X}$. Works such as scGen (Lotfollahi et al., 2019), CPA (Lotfollahi et al., 2023), ChemCPA (Hetzel et al., 2022), and CellOT (Bunne et al., 2023) aim to generalize to new settings $P_{X}$, e.g. unseen cell lines or patients. Alternatively, models like GEARS (Roohani et al., 2023), graphVCI (Wu et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) infer the effects of unseen perturbations. This setting is also a common application for single cell foundation models (Theodoris et al., 2023; Cui et al., 2024; Hao et al., 2024). These latter methods generalize by modeling the relationships between seen and unseen perturbations. For example, GEARS leverages literature-derived relationships in the Gene Ontology knowledge graph (Ashburner et al., 2000), and graphVCI learns to refine the gene regulatory relationships implied by ATAC-seq data (Grandi et al., 2022).
Approaches for inferring the effects of interventions do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2024). For example, given an oracle $\phi:D_{\text{obs}},I\to\tilde{P}$, the $I^{*}$ associated with $\tilde{P}^{*}$ is
@@
for some dissimilarity $d$. However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting that phenotype prediction is currently unsolved.
Perturbation datasets can also be used to train models for predicting targets of interest, as is the focus of this work. A common strategy is to relate observed changes (e.g. disease, gene levels) to mechanistic explanations (e.g. pathways), in which the targets are involved. To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2025). However, these information are collected from highly inhomogenous sources (Ashburner et al., 2000), and it is difficult to quantify their correctness and completeness in each new setting. An alternative is to infer the mechanistic structure directly from data, either explicitly (di Bernardo et al., 2005) or as a model parameter (Noh & Gunawan, 2016). In this work, we follow the latter strategy, while allowing for a richer (non-linear) class of relationships.
@@ 2.3 Causal discovery
If data from multiple environments are available, they can provide additional identifiability regarding the system (Jaber et al., 2020; Zhang et al., 2023b). Differences between these environments, e.g. intervention targets $I$, are either given as input or part of the inference task. As in the observational case, discovery algorithms can be categorized into discrete search (Hauser & Bühlmann, 2012; Ghassami et al., 2018; Ke et al., 2019; Huang et al., 2020; Mooij et al., 2020) or continuous optimization (Brouillard et al., 2020; Hägele et al., 2023). Additionally, Varici et al. (2022) and Yang et al. (2024) frame intervention target prediction as its own task, similar to this paper. However, the former strictly assumes linearity, while the latter requires at least as many environments as $\|I\|$ (we only consider before and after perturbation). Finally, several works aim to infer differences between causal mechanisms (edges), rather than identifying the full causal structure (Wang et al., 2018; Belyaeva et al., 2021; Chen et al., 2023; Malik et al., 2024). Our work is also motivated by the idea that differences between summary statistics (e.g. regression coefficients in Dci) are informative of changes between systems. Thus, Cdn can be viewed as implementing similar operations within an end-to-end, supervised framework.
@@ 2.4 Amortized causal inference
While the majority of causal discovery algorithms must be fit or run from scratch for each data distribution, amortized causal discovery algorithms have addressed this task as a supervised machine learning problem. Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2023; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2025). Compared to traditional approaches, these algorithms trade explicit assumptions for implicit characteristics of the data simulation process (Montagna et al., 2024). These models are also fast and robust to noise. In this work, we simulate interventions that reflect biological perturbations, and our model architecture builds upon that of Wu et al. (2025), as it scales easily to hundreds of variables.
@@ 3.1 Causal structure learner
The causal structure learner takes as input dataset $D$ and outputs the predicted $G$. In principle, this module can be implemented with any data-efficient causal discovery model. Here, we based the architecture on Sea (Wu et al., 2025), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
Input featurization  We represent dataset $D$ through two types of statistics. First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$ (pairwise correlation) over all variables. This statistic is an inexpensive way for the model to filter out edges that are not likely to occur. Second, we obtain estimates of local causal structure to orient edges, by running classical causal discovery algorithms over small subsets of variables. Specifically, we run the FCI algorithm $T$ times over subsets of $k$ variables (Spirtes et al., 1995), which are sampled proportional to their pairwise correlation. We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types inferred by FCI (e.g. ,$\leftrightarrow$,-$\rightarrow$ , - , →). The size and number of local estimates, $k$ and $T$, are hyperparameters that trade off accuracy vs. runtime (Section 3.3).
@@
where $h\in\mathbb{R}^{\dots\times d}$ denotes an intermediate hidden representation. This design takes after Transformer models for higher dimensional data in other domains, e.g. Axial Transformers for videos (Ho et al., 2019) and MSA Transformers for aligned protein sequences (Rao et al., 2021). A key difference is the graph modality: when making edge or node-level predictions, the model should be equivariant to node labeling. This is accomplished by adding randomly-permuted “positional” embeddings to each node and edge representation. For more details, please see Appendix A.
@@ 3.3 Implementation details
Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995) with $k=5$ and $T=100$, and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable. We swept over the number of differential network layers (Figure 4) on synthetic data, and we used 3 layers for $h_{\text{cat}}$ and 2 layers for $h_{\text{diff}}$. Following Sea, we adopted hidden dimension $d=64$, the AdamW optimizer (Loshchilov & Hutter, 2019), learning rate 1e-4, batch size 16, and weight decay 1e-5. On the real data, where $N=1000$, we changed to a batch size of 1, decreased the learning rate to 5e-6, and finetuned the models with half precision (FP16). We trained both $h_{\text{cat}}$ and $h_{\text{diff}}$ models on synthetic data, but only finetuned the $h_{\text{diff}}$ architecture on the real data (memory constraint).
@@ 4 Experiments
While our method is intended to be used in cases where true targets are unknown, we can only evaluate on datasets where the target is known. Thus, we demonstrate that Cdn can recover perturbation targets on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1). For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
@@ 4.1 Biological experiments
Datasets  We validate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al. (2022) and Nadig et al. (2024); as well as two Sci-Plex (Srivatsan et al., 2020) datasets (chemical perturbations) from McFaline-Figueroa et al. (2024). Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$. In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target (discussion in Appendix D.1).
@@
  • We order genes based on the similarity of their predicted effect to the ground truth target effect, or based on their predicted probability of being a target. The rank is the index of the true target, normalized by the number of candidate genes. Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
@@
  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates. Recall ranges from 0 to 1, where 1 is best. Rank is equivalent to the expected recall at $p$.
@@ 4.2 Synthetic experiments
exclude scale $0.32\scriptstyle\pm.07$ $0.64\scriptstyle\pm.04$ $0.34\scriptstyle\pm.09$ $0.65\scriptstyle\pm.08$ $0.06\scriptstyle\pm.02$ $0.63\scriptstyle\pm.02$ $0.08\scriptstyle\pm.02$ $0.61\scriptstyle\pm.06$
exclude scale 0.44±.04plus-or-minus0.44.040.$4$04 $0.92\scriptstyle\pm.01$ $0.43\scriptstyle\pm.09$ $0.90\scriptstyle\pm.03$ $0.06\scriptstyle\pm.01$ $0.72\scriptstyle\pm.02$ $0.10\scriptstyle\pm.02$ $0.73\scriptstyle\pm.02$
@@
5 Discussion
This work introduces causal differential networks (Cdn), which focuses on the application of predicting perturbation targets for efficient experimental design. On a high level, our goal is to lower the cost and scale of experiments, by reducing the search space. Our key technical contribution is the idea of using supervised causal discovery as a pretraining objective, to obtain dataset representations useful for downstream tasks like inferring targets. Specifically, Cdn trains a causal graph learner that maps a pair of observational and interventional datasets to their data-generating mechanisms, and a differential network classifier to identify variables whose conditional independencies have changed. We demonstrated that our approach outperforms state-of-the-art methods in perturbation modeling and causal discovery, across seven transcriptomics datasets and a variety of synthetic settings. Overall, these results highlight the potential of our method to design more efficient experiments, by reducing the search space of perturbations.
There are several limitations of this work and directions for future research. First, our implementation does not address cyclic, feedback relationships or time-resolved, dynamic systems – both of which are common in biology. Modeling these cases may require data simulation schemes that can sample from the steady-state of a cyclic system, or architectures that directly predict the parameters of a system of stochastic differential equations (Lorch et al., 2024). In addition, current supervised causal discovery algorithms, which we use for the causal graph learner, tend to assume causal sufficiency, i.e. that there are no unseen confounders. This assumption is unrealistic in biology, as it is not physically possible to measure all relevant variables. Thus, there is an opportunity to extend this work for the latent confounding setting. Finally, since it is difficult to simulate biology, there will be an inevitable domain shift between synthetic and real data. However, there does exist plentiful unlabeled single-cell data (Consortium* et al., 2022), which could be used with self-supervised objectives to bridge the gap (Sun et al., 2020). In conclusion, we hope that this work will enable efficient experimental design and interpretation, as well as future machine learning efforts towards these tasks.
@@ A.1 Attention layers
To balance expressivity and efficiency, our model is designed around series of “axial” attention layers, similar to Ho et al. (2019); Rao et al. (2021); Wu et al. (2025). The causal structure learner closely resembles the architecture of Sea (Wu et al., 2025), while the differential network is slightly different (since the input space is different). Briefly, the input to the causal structure learner is a matrix of shape $N\times N\times T\times d$. However, the (dense) global statistic does not span the $T$ dimension, and the local graph estimates are generally sparse. Thus, the causal graph learner separately attends over the $N\times N\times d$ and $K\times T\times d$ components ($K$ is the number of unique edges in sparse format), with “messages” to align the representations between each attention layer. In contrast, the differential network only operates over the dense $N\times(N+1)\times d$ (or $2d$) representation. Thus, it omits the message passing layers and directly stacks dense attention layers.